# MFCLIP: Multi-modal Fine-grained CLIP for Generalizable Diffusion Face Forgery Detection

### v1

@@ VI Visualization
Visualization of SPA. To demonstrate the effectiveness of our SPA method, we visualized the cosine similarity matrix, when SPA is involved or not. As Fig. 9 (a) displays, relevant negative pairs acquire larger correlation scores and vice versa, due to the addition of SPA, which shows that SPA emphasizes the related negative pairs and suppresses the irrelevant ones to achieve flexible alignment.
Visualization of heatmap. To further investigate the effect of MFCLIP, we displayed the heatmap of different detectors in Fig. 9 (b). Each row shows a forgery face yielded by various generators. The second to fifth columns illustrate heatmaps of four models: (I) CLIP; (II) DIRE; (III) FreqNet, and (IV) MFCLIP. Compared to other detectors, MFCLIP (III) captures more manipulated areas, showing that text-guided image-noise face forgery representation learning can benefit FFD. Specifically, in the last row, the pristine face is added with bangs through the attribute-manipulated model Diffae, and we noticed that the MFCLIP model could, to a large extent, identify the forgery area.
@@ VII Conclusion
Limitations. Although our model has explored fine-grained multi-modal forgery traces, we may need to improve generalization across diverse state-of-the-art datasets, and reduce computational complexity. In the future, we intend to integrate the pre-trained MVE with the multi-modal large language model, to improve the ability to understand and detect complex face forgeries.

### v2

@@ VI Visualization
Visualization of SPA. To demonstrate the effectiveness of our SPA method, we visualize the cosine similarity matrix, when SPA is involved or not. As Fig. 8 (a) displays, relevant negative pairs acquire larger correlation scores and vice versa, due to the addition of SPA, which shows that SPA emphasizes the related negative pairs and suppresses the irrelevant ones to achieve flexible alignment.
Visualization of heatmap. To further investigate the effect of MFCLIP, we display the heatmap of different detectors in Fig. 8 (b). Each row shows a forgery face yielded by various generators. The second to fifth columns illustrate heatmaps of four models: (I) CLIP; (II) DIRE; (III) FreqNet, and (IV) MFCLIP. Compared to other detectors, MFCLIP (III) captures more manipulated areas, showing that text-guided image-noise face forgery representation learning can benefit FFD. Specifically, in the last row, the pristine face is added with bangs through the attribute-manipulated model Diffae, and we notice that the MFCLIP model could, to a large extent, identify the forgery area.
@@ VII Conclusion
Limitations. Although our model has explored fine-grained multi-modal forgery traces, we need to improve generalization across various generators and reduce computational complexity. In the future, to solve the first restriction, fine-tuning the detector with face forgery data generated by advanced methods, such as autoregressive and flow-based models, could improve the generalizability and reliability. To overcome the second limitation, we are expected to devote ourselves to model architecture optimization such as lightweight model design and module replacement, for faster efficiency. For the former, reducing the number of transformer layers, hidden dimensions, or attention heads could be realized to decrease the complexity of the proposed model. In addition, a simpler and faster student model could be trained to mimic the behaviour of the proposed MFCLIP model via knowledge distillation. For the latter, linear attention or other efficient attention variants (e.g., vision mamba [62] or spike-driven self-attention [63]) could be employed to reduce the computational complexity of the transformer in MFCLIP.